# Computation of the expectation value of the spin operator $\hat{S}^{2}$ for the Spin-Flip Bethe-Salpeter Equation

### v1

@@ 6 Example application: The defect states of the ${}^{-}$ center
Figure 5: The electron configurations and values of $\langle\hat{S}^{2}\rangle$ as calculated in SF-BSE for (a) the high-spin reference state, (b) the ${}^{3}A_{2}$ (with $M_{S}=0$) ground state, and (c) the ${}^{3}E$ excited state for the ${}^{-}$ center in diamond. Values of $\langle\hat{S}^{2}\rangle$ are presented for both the spin-restricted and spin-unrestricted cases.
@@ 7 Execution and timing information of implementation with BerkeleyGW
In practice, the overlaps are calculated from files with potentially many more orbitals (or bands) than specified for the SF-BSE calculation; these total number of bands and their occupation values are then also needed. From this latter requirement, we read as input the wavefunction information calculated for the high-spin reference state, in which all unpaired electrons are in the spin-up channel. The wavefunction information, as currently implemented, is provided from an HDF5-format [34] file, wfn.h5, obtained through BerkeleyGW with HDF5 I/O enabled. In particular, the data “mnband”,“ifmin”, and “ifmax” from the wfn.h5 file are used. As per the BerkeleyGW documentation [35], “mnband”, the total number of bands, is a single integer. “ifmin” and “ifmax” are integer arrays with two indices; the first (in Python order – Fortran uses the reverse) is the spin index, and the second is the $k$-point index. With the possible exception of defects in two-dimensional materials [36, 37], only the $\Gamma$-point in the Brillouin zone is needed. “ifmin” is the lowest occupied state, while “ifmax” is the highest occupied state (resolved by spin and, if needed, $k$-point index). The combination of the data from “mnband”, “ifmin”, and “ifmax” is used to designate orbitals/bands in either spin channel as occupied or unoccupied.
@@
The output of this script should be piped to an output text file.
@@ 8 Conclusion
We showed the necessity of the computation of $\langle\hat{S}^{2}\rangle$ to make meaning of the results from Spin-Flip Bethe-Salpeter Equation calculations due to ambiguities from phase differences in orbitals in different spin channels, and ease of rapid identification of desired excited states. We reviewed the notion of spin contamination and its possible sources from spin-polarized calculations (generally giving small contamination) and incomplete transition vectors (possibly giving large contamination). We then derived in great detail the equation used for calculating $\Delta\langle\hat{S}^{2}\rangle$ from the “super-operator” approach, by explicitly considering the four cases possible for the individual transitions in the bra and ket states. We reviewed the suite of Python scripts written for BerkeleyGW to calculate $\langle\hat{S}^{2}\rangle_{0}$ for the high-spin reference state and $\Delta\langle\hat{S}^{2}\rangle$ for the excited states, so that this approach may be implemented independently for research teams interested in the spin-flip method. These scripts will be available in a future release of the BerkeleyGW package. We included timing information for the execution of the (serial) script for $\Delta\langle\hat{S}^{2}\rangle$, showing $O(N_{v}^{4}N_{c}^{3})$ scaling, when computing for all possible excitations. The convergence of $\Delta\langle\hat{S}^{2}\rangle$ for our two examples of ethylene and the ${}^{-}$ center in diamond indicated rapid convergence with respect to the number of unoccupied states, so the serial script with fewer unoccupied states may be performed initially for the purposes of analysis of states before the full set of unoccupied states are used for fully-converged results for $\Delta\langle\hat{S}^{2}\rangle$ values.

### v2

@@ 6 Example application: The defect states of the NV- center
Figure 5: The electron configurations and values of $\langle\hat{S}^{2}\rangle$ as calculated in SF-BSE for (a) the high-spin reference state, (b) the ${}^{3}A_{2}$ (with $M_{S}=0$) ground state, and (c) the ${}^{3}E$ excited state for the NV- center in diamond. Values of $\langle\hat{S}^{2}\rangle$ are presented for both the spin-restricted and spin-unrestricted cases.
@@ 7 Execution and timing information of implementation with BerkeleyGW
In practice, the overlaps are calculated from files with potentially many more orbitals (or bands) than specified for the SF-BSE calculation; these total number of bands and their occupation values are then also needed. From this latter requirement, we read as input the wavefunction information calculated for the high-spin reference state, in which all unpaired electrons are in the spin-up channel. The wavefunction information, as currently implemented, is provided from an HDF5-format [36] file, wfn.h5, obtained through BerkeleyGW with HDF5 I/O enabled. In particular, the data “mnband”,“ifmin”, and “ifmax” from the wfn.h5 file are used. As per the BerkeleyGW documentation [37], “mnband”, the total number of bands, is a single integer. “ifmin” and “ifmax” are integer arrays with two indices; the first (the fast index (first of Python/C, last for Fortran) is the spin index, and the second is the $k$-point index. With the possible exception of defects in two-dimensional materials [38, 39], only the $\Gamma$-point in the Brillouin zone is needed. “ifmin” is the lowest occupied state, while “ifmax” is the highest occupied state (resolved by spin and, if needed, $k$-point index). The combination of the data from “mnband”, “ifmin”, and “ifmax” is used to designate orbitals/bands in either spin channel as occupied or unoccupied.
@@
The output of this script should be piped to an output text file. The scripts to calculate $\langle\hat{S}^{2}\rangle$ are available publicly in Ref. [40].
@@ 8 Conclusion
We showed the necessity of the computation of $\langle\hat{S}^{2}\rangle$ to make meaning of the results from Spin-Flip Bethe-Salpeter Equation calculations due to ambiguities from phase differences in orbitals in different spin channels, and ease of rapid identification of desired excited states. We reviewed the notion of spin contamination and its possible sources from spin-polarized calculations (generally giving small contamination) and incomplete transition vectors (possibly giving large contamination). We then derived in great detail the equation used for calculating $\Delta\langle\hat{S}^{2}\rangle$ from the “super-operator” approach, by explicitly considering the four cases possible for the individual transitions in the bra and ket states. We reviewed the suite of Python scripts written for BerkeleyGW to calculate $\langle\hat{S}^{2}\rangle_{0}$ for the high-spin reference state and $\Delta\langle\hat{S}^{2}\rangle$ for the excited states, so that this approach may be implemented independently for research teams interested in the spin-flip method. These scripts will be available in a future release of the BerkeleyGW package. We included timing information for the execution of the (serial) script for $\Delta\langle\hat{S}^{2}\rangle$, showing $O(N_{v}^{4}N_{c}^{3})$ scaling, when computing for all possible excitations. The convergence of $\Delta\langle\hat{S}^{2}\rangle$ for our two examples of ethylene and the NV- center in diamond indicated rapid convergence with respect to the number of unoccupied states, so the serial script with fewer unoccupied states may be performed initially for the purposes of analysis of states before the full set of unoccupied states are used for fully-converged results for $\Delta\langle\hat{S}^{2}\rangle$ values.